OpenStudy (anonymous):

A new restaurant is to contain 2-seat tables and 4-seat tables. Fire codes limit the restaurant's maximum occupancy to 44 customers. If the owners have hired enough servers to handle 15 tables of customers, how many of each kind of table should they purchase?

OpenStudy (texaschic101):

t = two seat tables f = four seat tables t + f = 15 ---> t = 15 - f 2t + 4f = 44 now sub in 15 - f in for t in the 2nd equation 2(15 - f) + 4f = 44 30 - 2f + 4f = 44 --- subtract 30 from both sides -2f + 4f = 44 - 30 -- combine like terms 2f = 14 -- divide both sides by 2 f = 7 now sub 7 in for f in either equation to find t t + f = 15 t + 7 = 15 t = 15 - 7 t = 8 check.. 2t + 4f = 44 2(8) + 4(7) = 44 16 + 28 = 44 44 = 44 (correct) There are 7 four seat tables and 8 two seat tables

OpenStudy (anonymous):

You make it look so simple. Why do these function word questions always look harder than they are?

OpenStudy (anonymous):

Thankee-sai, btw.

OpenStudy (texaschic101):

sometimes they can be a bit tricky. You just got to make sure you read the question slowly and take your time....then check your answers. Do it enough and it will come easier to you :)

OpenStudy (anonymous):

Yeah, not so much. I can't make my brain understand the need for functions when every single thing I am studying does not use functions

OpenStudy (anonymous):

So, therefore, I can't make it stick :( Also, this is a distance learning class, so I am not able to listen to lecture and have that to reinforce the learning